# Entanglement and apparent thermality in simulated black holes

### v1

@@ 5 Bogoliubov transformation and Hawking effect
The Fulling-Davies-Unruh effect originates in the fact that the vacuum states of a Minkowski observer and a Rindler observer are not equivalent [4, 55]. An observer in a Minkowski space-time will define their vacuum state as the state $|0_{M}\rangle$ that satisfies $\hat{A}_{K}|0_{M}\rangle=\hat{B}_{K}|0_{M}\rangle=0$ $\forall K$, where $\hat{A}_{K}$ and $\hat{B}_{K}$ are the particle and anti-particle annihilation operators, respectively, associated with the general solution of Eq. (C.59). On the other hand, a Rindler observer will define their vacuum state as the state $|0_{R}\rangle$ that satisfies $\hat{a}_{k}|0_{R}\rangle=\hat{b}_{k}|0_{R}\rangle=0$ $\forall k$, where $\hat{a}_{k}$ and $\hat{b}_{k}$ are the particle and anti-particle annihilation operators, respectively, associated with the general solution of Eq. (C.40). Generally, the vacuum states $|0_{M}\rangle$ and $|0_{R}\rangle$ are not equivalent, as the quantum field theories for the Minkowski and Rindler observers have been constructed by quantising general solutions to the Dirac equation that were obtained for different coordinate systems. Nevertheless, the two quantum field theories should be equivalent when the Minkowski and Rindler coordinate systems are equivalent [64, 53]. This occurs at $\tilde{t}=T=0$, $\rho=X$; thus, at $\tilde{t}=T=0$, we have $\hat{\psi}(0,\rho)=\hat{\Psi}(0,X)$ and $\hat{\pi}(0,\rho)=\hat{\Pi}(0,X)$ [53]. As the creation and annihilation operators are determined by each theory’s Cauchy initial data $\hat{\psi}(0,\rho)$ and $\hat{\Psi}(0,X)$, their equivalence at $\tilde{t}=T=0$ allows each theory’s creation and annihilation operators to be expressed in terms of the other’s. That is,
$\begin{split}\hat{a}_{k}&=\langle\hat{\psi}_{k},\hat{\psi}(0,\rho)\rangle\\ &=\int d\mu(K)\left(\hat{A}_{K}\langle\hat{\psi}_{k},\hat{\Psi}_{K}\rangle+\hat{B}_{K}^{\dagger}\langle\hat{\psi}_{k},\hat{\Phi}_{K}\rangle\right)\end{split}$ (C.60)
@@
$\begin{split}\hat{b}^{\dagger}_{k}&=\langle\hat{\phi}_{k},\hat{\psi}(0,\rho)\rangle\\ &=\int d\mu(K)\left(\hat{A}_{K}\langle\hat{\phi}_{k},\hat{\Psi}_{K}\rangle+\hat{B}_{K}^{\dagger}\langle\hat{\phi}_{k},\hat{\Phi}_{K}\rangle\right),\end{split}$ (C.61)
where the equivalence of the two integration measures $d\mu(\rho)$ and $d\mu(X)$ at $\tilde{t}=T=0$ has been used to compute the inner product between the Minkowski and Rindler eigenstates. Eqs. (C.60) and (C.61) each give a Bogoliubov transformation between the Minkowski and Rindler creation and annihilation operators [51, 49, 53]. Using these Bogoliubov transformations and the fact that $\hat{A}_{K}|0_{M}\rangle=\langle 0_{M}|\hat{A}^{\dagger}_{K}=0$, the vacuum expectation value of the Minkowski vacuum, as seen by the Rindler observer, can be expressed as
$\langle 0_{M}|\hat{a}^{\dagger}_{k}\hat{a}_{q}|0_{M}\rangle=\int d\mu(K)\langle\hat{\phi}_{k},\hat{\Psi}_{K}\rangle\langle\hat{\psi}_{q},\hat{\Phi}_{K}\rangle,$ (C.62)
@@
$=\frac{u^{T}_{q}u_{K}}{2\pi\sqrt{|\kappa|}}\int^{\infty}_{0}dX\left(X^{-\frac{iq}{\kappa}-\frac{1}{2}}e^{-iKX}\right),$ (C.64)
$=\frac{\delta_{qK}}{2\pi\sqrt{|\kappa|}}e^{-\frac{\pi q}{2\kappa}-\frac{i\pi}{4}}K^{\frac{iq}{\kappa}-\frac{1}{2}}\Gamma\left(\frac{1}{2}-\frac{iq}{\kappa}\right),$ (C.65)
@@
where $z=iKX$, $\Gamma(z)$ is the gamma function, and, in the final equality, the identity $i=e^{\frac{i\pi}{2}}$ was used;
$\langle\hat{\phi}_{k},\hat{\Psi}_{K}\rangle$ $=\int^{\infty}_{0}d\mu(X)\hat{\phi}^{\dagger}_{k}(X)\hat{\Phi}_{K}(X),$ (C.66)
@@
$=\frac{u^{T}_{k}u_{K}}{2\pi\sqrt{|\kappa|}}\int^{\infty}_{0}dX\left(X^{\frac{ik}{\kappa}-\frac{1}{2}}e^{iXK}\right),$ (C.67)
$=\frac{\delta_{kK}}{2\pi\sqrt{|\kappa|}}e^{-\frac{\pi k}{2\kappa}+\frac{i\pi}{4}}K^{-\frac{ik}{\kappa}+\frac{1}{2}}\Gamma^{*}\left(\frac{1}{2}-\frac{ik}{\kappa}\right),$ (C.68)
@@
where z=iKXz=-iKX, $\Gamma^{*}(z)=\Gamma(z^{*})$, and, in the final equality, the identity $i=e^{-\frac{i\pi}{2}}$ was used. Hence, the vacuum expectation value of the Minkowski vacuum, as seen by the Rindler observer, can be found to be
@@
$\begin{split}\langle 0_{M}|\hat{a}^{\dagger}_{k}\hat{a}_{q}|0_{M}\rangle&=\frac{\delta_{kq}}{2\pi}e^{-\frac{\pi}{2\kappa}(k+q)}\Gamma^{*}\left(\frac{1}{2}-\frac{ik}{\kappa}\right)\\ &\ \ \ \ \ \Gamma\left(\frac{1}{2}-\frac{iq}{\kappa}\right)\delta(k-q),\end{split}$ (C.69)
@@
$\langle 0_{M}|\hat{a}^{\dagger}_{k}\hat{a}_{k}|0_{M}\rangle$ $=\frac{\delta(0)}{2}e^{-\frac{\pi k}{\kappa}}\text{sech}\left(\frac{\pi k}{\kappa}\right),$ (C.70)
$=\frac{\delta(0)}{e^{2\pi k/\kappa}+1},$ (C.71)
@@
where, in the second equality, the identity $|\Gamma(ix+\frac{1}{2})|^{2}=\pi\text{sech}(\pi x)$ has been used. Eq. (C.71) for the Minkowski vacuum expectation value, as seen by the Rindler observer, is a Fermi-Dirac distribution for fermions in thermal equilibrium with a temperature $T_{H}=\kappa/2\pi$, where $\kappa=-v^{\prime}(x_{h})$. That is, an observer that is accelerating with a constant proper acceleration $\kappa$ will experience their vacuum state to be populated by thermal radiation with a temperature $T_{H}$ [53, 54, 55]; this is the Fulling-Davies-Unruh effect. As the Schwarzschild metric can be approximated by the Rindler metric near the event horizon, an observer near the event horizon of a Schwarzschild black hole will experience the Fulling-Davies-Unruh effect [4, 61].
@@
Appendix D. Fermion zero-mode entropy of (1+1)D black hole exterior
@@
In this appendix, we extend the analysis of Sec. III to derive an expression for the fermion zero-mode entropy of both the black hole’s interior and exterior regions. The method of this derivation is analogous to that of Sec. III and involves determining the density of states of Eq. (13) for both the interior region, which is bound by $0,x_{h}-a_{c}]$, and a portion of the exterior space-time, which we take to be bound by $x_{h}+a_{c},s]$, where $s\geq x_{h}+a_{c}$. To prevent an infinite number of degrees of freedom, from points on either side of the event horizon with a separation of less than $a_{c}$, from contributing to the density of states, the lower limit of $s$ has been restricted to a distance $a_{c}$ from the event horizon [43, 44, 33].
From Eq. (13), the density of states for both the interior and exterior regions will be given by
$\begin{split}N(0)&=\frac{N_{F}}{\pi\hbar}\int^{x_{h}-a_{c}}_{0}dx\left(\frac{1}{|v(x)|}+\frac{1}{2}\Biggl|\frac{1}{\left(1-v(x)\right)}-\frac{1}{\left(1+v(x)\right)}\Biggr|^{-1}\right)\\ &\ \ \ \ \ \ \ \ \ +\frac{\tilde{N}_{F}}{\pi\hbar}\int^{s}_{x_{h}+a_{c}}dx\left(\frac{1}{|v(x)|}+\frac{1}{2}\Biggl|\frac{1}{\left(1-v(x)\right)}-\frac{1}{\left(1+v(x)\right)}\Biggr|^{-1}\right),\end{split}$ (D.1)
@@
where $\tilde{N}_{F}$ is the number of fermion species in the exterior region of the black hole. Taking the coupling $v(x)=\sqrt{x_{h}/x}$, which corresponds to that of a Schwarzschild black hole, and working in the limit $0<a_{c}\ll x_{h}$, this density of states can be solved to yield
$N(0)=\frac{N_{F}x_{h}}{\pi\hbar}\left(\ln\left(\frac{4x_{h}}{a_{c}}\right)-\frac{4}{3}\right)+\frac{\tilde{N}_{F}x_{h}}{\pi\hbar}\left(\frac{2}{3}\left(\frac{s}{x_{h}}\right)^{\frac{3}{2}}+2\left(\frac{s}{x_{h}}\right)^{\frac{1}{2}}+\ln\left|\frac{{s}^{\frac{1}{2}}-x_{h}^{\frac{1}{2}}}{{s}^{\frac{1}{2}}+x_{h}^{\frac{1}{2}}}\right|+\ln\left(\frac{4x_{h}}{a_{c}}\right)-\frac{8}{3}\right),$ (D.2)
@@
where the term preceded by $N_{F}$ corresponds to the density of states of the entire black hole interior, and the term preceded by $\tilde{N}_{F}$ corresponds to the density of states of the black hole exterior.
@@
Assuming that the fermions in the black hole’s interior region are in thermal equilibrium with the exterior, such that their thermal energy and, hence, entropy are given by Eq. (16) and (17), respectively, the fermion zero-mode entropy of the black hole’s interior and exterior regions will be given by
$\mathcal{S}(T)=\frac{\pi N_{F}x_{h}}{3\hbar}\left(\ln\left(\frac{4x_{h}}{a_{c}}\right)-\frac{4}{3}\right)T+\frac{\pi\tilde{N}_{F}x_{h}}{3\hbar}\left(\frac{2}{3}\left(\frac{s}{x_{h}}\right)^{\frac{3}{2}}+2\left(\frac{s}{x_{h}}\right)^{\frac{1}{2}}+\ln\left|\frac{{s}^{\frac{1}{2}}-x_{h}^{\frac{1}{2}}}{{s}^{\frac{1}{2}}+x_{h}^{\frac{1}{2}}}\right|+\ln\left(\frac{4x_{h}}{a_{c}}\right)-\frac{8}{3}\right)T,$ (D.3)
where $T$ is the temperature at which the fermions are in equilibrium. Once again, assuming that this temperature is the Hawking temperature of Eq. (18), the entropy of the fermion zero-modes in the black hole’s interior and exterior will be given by
$\mathcal{S}(T_{H})=\frac{N_{F}}{12}\ln\left(\frac{4x_{h}}{a_{c}}\right)+\frac{\tilde{N}_{F}}{12}\left(\frac{2}{3}\left(\frac{s}{x_{h}}\right)^{\frac{3}{2}}+2\left(\frac{s}{x_{h}}\right)^{\frac{1}{2}}+\ln\left|\frac{{s}^{\frac{1}{2}}-x_{h}^{\frac{1}{2}}}{{s}^{\frac{1}{2}}+x_{h}^{\frac{1}{2}}}\right|+\ln\left(\frac{4x_{h}}{a_{c}}\right)\right)+\mathcal{S}_{0},$ (D.4)
@@
where $\mathcal{S}_{0}=-(N_{F}-2\tilde{N}_{F})/9$. For the case where $s=x_{h}+a_{c}$, and in the limit $a_{c}\ll x_{h}$, Eq. (D.4) for the fermion zero-mode entropy of the black hole’s interior and exterior reduces to Eq. (20), as expected.

### v2

@@ 5 Bogoliubov transformation and Hawking effect
The Fulling-Davies-Unruh effect originates in the fact that the vacuum states of a Minkowski observer and a Rindler observer are not equivalent [69, 60]. An observer in a Minkowski space-time will define their vacuum state as the state $|0_{M}\rangle$ that satisfies $\hat{A}_{K}|0_{M}\rangle=\hat{B}_{K}|0_{M}\rangle=0$ $\forall K$, where $\hat{A}_{K}$ and $\hat{B}_{K}$ are the particle and anti-particle annihilation operators, respectively, associated with the general solution of Eq. (B.52). On the other hand, a Rindler observer will define their vacuum state as the state $|0_{R}\rangle$ that satisfies $\hat{a}_{k}|0_{R}\rangle=\hat{b}_{k}|0_{R}\rangle=0$ $\forall k$, where $\hat{a}_{k}$ and $\hat{b}_{k}$ are the particle and anti-particle annihilation operators, respectively, associated with the general solution of Eq. (B.36). Generally, the vacuum states $|0_{M}\rangle$ and $|0_{R}\rangle$ are not equivalent, as the quantum field theories for the Minkowski and Rindler observers have been constructed by quantising general solutions to the Dirac equation that were obtained for different coordinate systems. Nevertheless, the two quantum field theories should be equivalent when the Minkowski and Rindler coordinate systems are equivalent [25, 23]. This occurs at $\tilde{t}=T=0$, $\rho=X$; thus, at $\tilde{t}=T=0$, we have $\hat{\psi}(0,\rho)=\hat{\Psi}(0,X)$ and $\hat{\pi}(0,\rho)=\hat{\Pi}(0,X)$ [23]. As the creation and annihilation operators are determined by each theory’s Cauchy initial data $\hat{\psi}(0,\rho)$ and $\hat{\Psi}(0,X)$, their equivalence at $\tilde{t}=T=0$ allows each theory’s creation and annihilation operators to be expressed in terms of the other’s. That is,
$\begin{split}\hat{a}_{k}&=\langle\hat{\psi}_{k},\hat{\psi}(0,\rho)\rangle\\ &=\int d\mu(K)\left(\hat{A}_{K}\langle\hat{\psi}_{k},\hat{\Psi}_{K}\rangle+\hat{B}_{K}^{\dagger}\langle\hat{\psi}_{k},\hat{\Phi}_{K}\rangle\right)\end{split}$ (B.53)
@@
$\begin{split}\hat{b}^{\dagger}_{k}&=\langle\hat{\phi}_{k},\hat{\psi}(0,\rho)\rangle\\ &=\int d\mu(K)\left(\hat{A}_{K}\langle\hat{\phi}_{k},\hat{\Psi}_{K}\rangle+\hat{B}_{K}^{\dagger}\langle\hat{\phi}_{k},\hat{\Phi}_{K}\rangle\right),\end{split}$ (B.54)
where the equivalence of the two integration measures $d\mu(\rho)$ and $d\mu(X)$ at $\tilde{t}=T=0$ is used to compute the inner product between the Minkowski and Rindler eigenstates. Eqs. (B.53) and (B.54) each give a Bogoliubov transformation between the Minkowski and Rindler creation and annihilation operators [20, 41, 23]. Using these Bogoliubov transformations and the fact that $\hat{A}_{K}|0_{M}\rangle=\langle 0_{M}|\hat{A}^{\dagger}_{K}=0$, the vacuum expectation value of the Minkowski vacuum, as seen by the Rindler observer, can be expressed as
$\langle 0_{M}|\hat{a}^{\dagger}_{k}\hat{a}_{q}|0_{M}\rangle=\int d\mu(K)\langle\hat{\phi}_{k},\hat{\Psi}_{K}\rangle\langle\hat{\psi}_{q},\hat{\Phi}_{K}\rangle,$ (B.55)
@@
$\langle\hat{\psi}_{q},\hat{\Phi}_{K}\rangle$ $=\frac{u^{T}_{q}u_{K}}{2\pi\sqrt{|\kappa|}}\int^{\infty}_{0}dX\left(X^{-\frac{iq}{\kappa}-\frac{1}{2}}e^{-iKX}\right),$ (B.56)
$=\frac{\delta_{qK}}{2\pi\sqrt{|\kappa|}}e^{-\frac{\pi q}{2\kappa}-\frac{i\pi}{4}}K^{\frac{iq}{\kappa}-\frac{1}{2}}\Gamma\left(\frac{1}{2}-\frac{iq}{\kappa}\right),$ (B.57)
@@
where denotes the $\Gamma(z)$ is the gamma function and, in the final equality, the identity $i=e^{\frac{i\pi}{2}}$ was used;
@@
$\langle\hat{\phi}_{k},\hat{\Psi}_{K}\rangle$ $=\frac{u^{T}_{k}u_{K}}{2\pi\sqrt{|\kappa|}}\int^{\infty}_{0}dX\left(X^{\frac{ik}{\kappa}-\frac{1}{2}}e^{iXK}\right),$ (B.58)
$=\frac{\delta_{kK}}{2\pi\sqrt{|\kappa|}}e^{-\frac{\pi k}{2\kappa}+\frac{i\pi}{4}}K^{-\frac{ik}{\kappa}+\frac{1}{2}}\Gamma^{*}\left(\frac{1}{2}-\frac{ik}{\kappa}\right),$ (B.59)
@@
where $\Gamma^{*}(z)=\Gamma(z^{*})$ and, in the final equality, the identity $i=e^{-\frac{i\pi}{2}}$ was used. Hence, the vacuum expectation value of the Minkowski vacuum, as seen by the Rindler observer, can be found to be
@@
$\begin{split}\langle 0_{M}|\hat{a}^{\dagger}_{k}\hat{a}_{q}|0_{M}\rangle&=\frac{\delta_{kq}}{2\pi}e^{-\frac{\pi}{2\kappa}(k+q)}\Gamma^{*}\left(\frac{1}{2}-\frac{ik}{\kappa}\right)\\ &\ \ \ \ \ \ \Gamma\left(\frac{1}{2}-\frac{iq}{\kappa}\right)\delta(k-q),\end{split}$ (B.60)
@@
$\langle 0_{M}|\hat{a}^{\dagger}_{k}\hat{a}_{k}|0_{M}\rangle$ $=\frac{\delta(0)}{2}e^{-\frac{\pi k}{\kappa}}\text{sech}\left(\frac{\pi k}{\kappa}\right),$ (B.61)
$=\frac{\delta(0)}{e^{2\pi k/\kappa}+1},$ (B.62)
@@
where, in the second equality, the identity $|\Gamma(ix+\frac{1}{2})|^{2}=\pi\text{sech}(\pi x)$ has been used. Eq. (B.62) for the Minkowski vacuum expectation value, as seen by the Rindler observer, is a Fermi-Dirac distribution for fermions in thermal equilibrium with a temperature $T_{H}=\kappa/2\pi$. That is, an observer accelerating with a constant proper acceleration $\kappa$ will experience their vacuum state to be populated by thermal radiation with a temperature $T_{H}$ [23, 19, 60]; this is the Fulling-Davies-Unruh effect. As the Schwarzschild metric can be approximated by the Rindler metric near the event horizon, an observer near the event horizon of a Schwarzschild black hole will experience the Fulling-Davies-Unruh effect [69, 68].
@@
Appendix C. Fermion zero-mode entropy of (1+1)D black hole exterior
@@
In this appendix, we extend the analysis of Sec. III to derive the fermion zero-mode entropy of both the black hole’s interior and exterior regions. This derivation is analogous to that of Sec. III and involves determining the density of states of Eq. (12) for both the interior region, which is bound by $a_{c},x_{h}-a_{c}]$, and a portion of the exterior space-time, which we take to be bound by $x_{h}+a_{c},s]$, where $s\geq x_{h}+a_{c}$. To prevent an infinite number of degrees of freedom, from points on either side of the event horizon with a separation of less than $a_{c}$, from contributing to the density of states, the lower limit of $s$ has been restricted to a distance $a_{c}$ from the event horizon [7, 9, 31].
Taking the sub-luminal dispersion relation of Eq. (5), and considering the emergence of the two zero-modes at momenta $p_{\pm}$ in the interior region, along with the zero-mode at $p_{0}$ that exists throughout the entire space-time, the density of states of Eq. (11) for the interior and exterior regions will be given by
$N(0)=\frac{N_{F}}{2\pi\hbar}\int^{x_{h}-a_{c}}_{a_{c}}dx\left(\frac{1}{|1-v(x)|}+\Biggl|\frac{1}{\left(1-v(x)\right)}-\frac{1}{\left(1+v(x)\right)}\Biggr|\right)+\frac{\tilde{N}_{F}}{2\pi\hbar}\int^{s}_{x_{h}+a_{c}}dx\left(\frac{1}{|1-v(x)|}\right),$ (C.1)
@@
where $\tilde{N}_{F}$ denotes the integer number of massless fermionic fields in the exterior region. Taking the coupling $v(x)=\sqrt{x_{h}/x}$ that corresponds to the space-time of a Schwarzschild black hole, and working in the limit $0<a_{c}\ll x_{h}$, this density of states can be solved to yield
$N(0)=\frac{N_{F}x_{h}}{\pi\hbar}\left(2\ln\left(\frac{2x_{h}}{a_{c}}\right)+\ln(2)-\frac{7}{2}\right)+\frac{\tilde{N}_{F}x_{h}}{\pi\hbar}\left(\frac{1}{2}\left(\frac{s}{x_{h}}\right)+\left(\frac{s}{x_{h}}\right)^{\frac{1}{2}}+\ln\left|{s}^{\frac{1}{2}}-x_{h}^{\frac{1}{2}}\right|+\ln\left(\frac{2x_{h}^{\frac{1}{2}}}{a_{c}}\right)-\frac{3}{2}\right),$ (C.2)
@@
where the term preceded by $N_{F}$ corresponds to the density of states of the entire black hole interior, and the term preceded by $\tilde{N}_{F}$ the density of states of the black hole exterior.
@@
Assuming that the fermions in the black hole’s interior and exterior regions are well described by a thermal Gibbs state, their thermal energy and entropy can be defined by Eqs. (15) and (16), respectively. Once again, further assuming that the temperature associated with this thermal state is the Hawking temperature of Eq. (17), the fermion zero-mode entropy of the black hole’s interior and exterior will be given by
$\mathcal{S}(T_{H})=\frac{N_{F}}{6}\ln\left(\frac{2x_{h}}{a_{c}}\right)+\frac{\tilde{N}_{F}}{12}\left(\frac{1}{2}\left(\frac{s}{x_{h}}\right)+\left(\frac{s}{x_{h}}\right)^{\frac{1}{2}}+\ln\left(\frac{2x_{h}^{\frac{1}{2}}}{a_{c}}\right)+\ln\left|{s}^{\frac{1}{2}}-x_{h}^{\frac{1}{2}}\right|\right)+\mathcal{S}^{\prime}_{0}$ (C.3)
@@
where $\mathcal{S}^{\prime}_{0}=\mathcal{S}_{0}-3\tilde{N}_{F}/24$. For the case where $s=x_{h}+a_{c}$, and in the limit $a_{c}\ll x_{h}$, Eq. (C.3) for the fermion zero-mode entropy of the black hole’s interior and exterior reduces to Eq. (19), as expected.
Appendix D. Conformal Invariance of Lattice Simulator
In this appendix, we study the correspondence between a transformation of the mean-field Hamiltonian and a change in coordinates of the continuum limit Hamiltonian. This correspondence arises due to an equivalence between the choice of spinor renormalisation and a conformal scaling of the emergent metric. In the continuum limit, this transformation preserves the curvature of the underlying space-time; thus, scalar invariant quantities remain unchanged.
In deriving the continuum limit of the mean-field Hamiltonian of Eq. (2), the choice of zweibein defined by Eq. (9) is not unique. One could alternatively define another set of coefficients $\tilde{e}^{\mu}_{a}$ that are related to $e^{\mu}_{a}$ via the transformation $e^{\mu}_{a}=C(u,v)\tilde{e}^{\mu}_{a}$, where, as we will see, $C(u,v)$ is a conformal scale factor. Following this transformation, the mean-field continuum limit Hamiltonian of Eq. (7) can be expressed as
$H=\int dp\penalty 10000\ C(u,v)\chi^{\dagger}(p)\tilde{e}^{x}_{a}\alpha^{a}p\chi(p),$ (D.4)
which, following an inverse Fourier transformation and, subsequently, a Legendre transformation, can be used to obtain the action
$S=i\int d^{1+1}x|\tilde{e}|\bar{\varphi}(x)\tilde{e}^{\mu}_{a}\gamma^{a}\overset{\leftrightarrow}{\partial}_{\mu}\varphi(x).$ (D.5)
Here, the curved space-time spinor $\varphi$, which is related to the lattice spinor $\chi$ via the renormalisation $\chi=\sqrt{|\tilde{e}|/C}\varphi$ and satisfies the anti-commutation relations of Eq. (A.20), and its conjugate $\bar{\varphi}=\varphi^{\dagger}(x)\gamma^{0}$, have been defined.
The action of Eq. (D.5) is, once again, that of a massless Dirac spinor $\varphi$ propagating freely on a (1+1)-dimensional Riemann-Cartan space-time, with the zweibein $\tilde{e}^{\mu}_{a}$ corresponding to the space-time metric line element
$d\tilde{s}^{2}=C(u,v)^{2}\left[\left(1-\frac{v^{2}}{u^{2}}\right)dt^{2}-\frac{2v}{u^{2}}dtdx-\frac{dx^{2}}{u^{2}}\right].$ (D.6)
From the line element of Eq. (D.6), we see that $C(u,v)$ is a conformal scale factor that relates the Gullstrand-Painlevé line element of Eq. (10) to that of Eq. (D.6) via the conformal transformation $d\tilde{s}^{2}=C(u,v)^{2}ds^{2}$.
A unique case of the general curved space-time of Eq. (D.6) can be obtained by taking $C(u)=\sqrt{u(x)}$ and working in the limit where $v\rightarrow 0$. In this case, the line element of Eq. (D.6) reduces to
$d\tilde{s}^{2}=u(x)dt^{2}-\frac{dx^{2}}{u(x)},$ (D.7)
which is that of the (1+1)-dimensional Schwarzschild metric. In this case, the corresponding lattice Hamiltonian is that of Eq. (2) with a position-dependent nearest-neighbour coupling $u(x)$ and vanishing chirality term.
To illustrate the invariance of the space-time’s curvature under this conformal transformation, it is instructive to study the Ricci curvature scalar $R$, defined as $R=g^{\mu\nu}R_{\mu\nu}$, where $R_{\mu\nu}=R^{\lambda}_{\mu\lambda\nu}$ denotes the Ricci tensor, which is itself defined in terms of the Riemann tensor $R_{\alpha\mu\lambda\nu}$. In a local inertial reference frame, the Riemann tensor is given by [10]
$R_{\alpha\mu\lambda\nu}=\frac{1}{2}\left(g_{\alpha\nu,\mu\lambda}-g_{\alpha\lambda,\mu\nu}+g_{\mu\lambda,\alpha\nu}-g_{\mu\nu,\alpha\lambda}\right),$ (D.8)
thus, for a stationary (1+1)-dimensional space-time metric, the components of the Ricci tensor can be found to be $R_{tt}=-\frac{1}{2}g^{xx}g_{tt,xx}$, $R_{xt}=R_{tx}=\frac{1}{2}g^{xt}g_{tt,xx}$, and $R_{xx}=-\frac{1}{2}g^{tt}g_{tt,xx}$. For the space-time metric of Eq. (D.6), the Ricci scalar is, therefore, given by
$R=\frac{u(x)^{2}}{C(x)^{4}}\frac{d^{2}}{dx^{2}}\left[C(x)^{2}\left(1-\frac{v(x)^{2}}{u(x)^{2}}\right)\right].$ (D.9)
It can be verified from the Ricci scalar of Eq. (D.9) that the curvature remains invariant under the conformal transformation and renormalisation of the curved space-time spinor, provided that the couplings are transformed such that they correspond to the same space-time geometry. Take, for example, the limit where the line element of Eq. (D.6) reduces to the Gullstrand-Painlevé line element of Eq. (10). In this limit, and for the couplings corresponding to the geometry of a (1+1)-dimensional Schwarzschild black hole, $u=1$ and $v(x)=\sqrt{x_{h}/x}$, the Ricci scalar is $R=-2x_{h}/x^{3}$. This is equivalent to the value obtained in the limit where the line element of Eq. (D.6) reduces to the Schwarzschild metric of Eq. (D.7) with the couplings of the (1+1)-dimensional Schwarzschild black hole geometry, $u(x)=(1-x_{h}/x)$ and $v=0$.